OpenStudy (anonymous):

For the equation f(x)=3/2(x+1)(x-4), how do I find the points in order to graph it?

OpenStudy (bgrg007):

u find the points where the graph meets the x axis and that would be x+1 = 0 , x=-1 x-4=0 , x=4

OpenStudy (anonymous):

Well in theory, by substituting random x-values, you can generate a table of values. And then plot it. Also, in theory, the more points you have, and the closer the points are, the better approximation of a curve you would obtain.

OpenStudy (anonymous):

To continue that point, as the number of points you have approach \(\infty\), and the space between the points approach \(\frac{1}{\infty}\), the approximation will approximate the function the best. Of course, the line joining infinite generated points adhering to \(f(x)\) would be \(f(x)\) itself!

OpenStudy (anonymous):

Is this a calculus question?

OpenStudy (anonymous):

But say I'm looking for whole number values for the y-coordinates, how do I know which x-values to substitute in?

OpenStudy (anonymous):

Advanced functions